Suppose your company has excess case of $13B, debt of $10B and 1 billion shares outstanding with a current price of $31 per share. ASsume its equity beta is 1.4

A. What is the enterprise value?

B. What is the current D/E ratio?

C. What is the asset (unlevered) beta?
Answer #1

Answer A)

Entereprise Value = Equity Value + Debt Value - Cash

= 1* 31 + 10 - 13

= 28 Bn

Answer B)

Debt Equity Ratio = Debt Value / Equity Value

= 10/ 31

= 0.32

Answer C)

Unlevered beta = Equity Beta 1- Equity

= 1.4/ ( 1- 0.32)

= 2.0588